OpenStudy (imtant):

How to integrate e^x^2 * x^3?

OpenStudy (anonymous):

This?\[ \int \exp(x^2)x^3\;dx \]

OpenStudy (anonymous):

\[ \int \exp(x^2)x^3\;dx=\int \exp(x^2)x^2\;(xdx) \]Let \(u=x^2,du=2xdx\)\[ \frac{1}{2}\int ue^u\;du \]

OpenStudy (anonymous):

\(\exp(x)=e^x\)

OpenStudy (imtant):

And x to the power of 2.

OpenStudy (anonymous):

\[ \exp(x^2)=e^{x^2} \]

OpenStudy (imtant):

Yes,that's it.

OpenStudy (anonymous):

\[ \exp(x^2)x^3=e^{x^2}x^3 \]See? It's the same thing dude.

OpenStudy (imtant):

Oh sorry.But I don't learn like that before.

OpenStudy (anonymous):

It's better not to make towers \[\large e^{x^{x^{e^{x^{x}}}}} \]

OpenStudy (imtant):

Okay.

OpenStudy (imtant):

But how about x^3?

OpenStudy (imtant):

∫exp(x2)x3dx=∫exp(x2)x2(xdx) Let u=x2,du=2xdx 12∫ueudu So if there is 2∫exp(x2)x3dx this will become 2*12∫ueudu, right?